# Integrating Multi-Physics Simulations and Machine Learning to Define the Spatter Mechanism and Process Window in Laser Powder Bed Fusion

### v1

@@ 4 Conclusion
In this work, we developed multiphysics simulation of a single-track selective laser melting using OpenFOAM, which has the capability to produce the 3D high resolution properties of spatter and meltpool dynamics. In addition, tracking algorithm was developed to extract the dataset used to characterize spatter ejection, the meltpool, and to perform ML classification tasks. Upon screening our datasets using different machine learning models such as ExtraTrees, XGB, LGBM, RF, Bagging, SVC and KNN, we observed high accuracy greater than 93 % for all the ML models implemented with the highest being ExtraTrees and the lowest being KNN. Upon removing spatial component (x, y, z) from the OpenFOAM dataset used to train/test ML model so as to match the accessible dataset that can be extracted from FLOW-3D, we found ExtraTrees and Random Forest to exhibit highest values at 93 % while KNN has the lowest value of 86 %. Random Forest shows least changes of 0.9 % when compared to their respective initial accuracy that was achieved with spatial component being included in the features. High accuracy outcome of ML predictions underscore the efficacy of our ML models in exploring the model for the reduced-order task of spatter prediction, by rigorously testing datasets from the commercial software. Region association with spatter on FLOW-3D software was predicted and identified as voxel. Simulating additive manufacturing process is 100 times faster on FLOW-3D than OpenFOAM, making screening of 281 different process parameters collected from FLOW-3D a feasible path towards developing a process map using ML model trained on OpenFOAM dataset a means of displaying spatter trend in large process parameter space. Correlation of the possible amount of spatter with changing power/velocity combination was generated. It was observed that at a high power and low laser speed, more spatters were generated in comparison to low power and high speed.
@@
Research was sponsored by the Army Research Laboratory and was accomplished under Cooperative Agreement Number W911NF-20-2-0175. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Laboratory or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@
Appendix A OpenFOAM simulation domain
Figure 10: Schematic illustration of the computational domain with random packing of Stainless steel powder bed
Table 3: Physical properties and related coefficients of SST in the simulation with subscript s for solid phase, l for liquid, v for vapor and a for the shielding gas argon.
Parameter Value Units
Liquidus temperature, $\rho$, 298 K 7950 kg/m3
Evaporation Temperature, $\rho$, 1923 K 6765 kg/m3
Surface tension liquid, $C_{v}$, 298 K 470 J/kg/K
Specific heat solid, $C_{v}$, 1923 K 1873 J/kg/K
Specific heat liquid, $C_{v,vapor}$ 449 J/kg/K
Specific heat vapor, $k$, 298 K 13.4 W/m/K
Specific heat nitrogen, $k$, 1923 K 30.5 W/m/K
Thermal conductivity solid, $\eta$ 0.008 kg/m/s
Thermal conductivity liquid, $\sigma$ 1.882 $kg/s^{2}$
Heat of fusion, $T_{L}$ 1723 K
Heat of evaporation, $T_{S}$ 1658 K
Dynamic viscosity liquid, $\epsilon$ 0.15 -
Dynamic viscosity vapor, $a$ 0.25 -
Prandtl number liquid, $\Delta H_{f}$ 2.6 ×$\times 10^{5}$ J/kg
Prandtl number vapor, $\Delta H_{v}$ 7.45 ×$\times 10^{6}$ J/kg
Absorptivity, $a$ 0.25 -

### v2

@@ 4 Conclusion
In this work, we developed multiphysics simulation of a single-track selective laser melting using OpenFOAM, which has the capability to produce the 3D high resolution properties of spatter and melt pool dynamics. In addition, a tracking algorithm was developed to extract the data set used to characterize spatter ejection, the melt pool, and to perform ML classification tasks. Upon screening our datasets using different machine learning models such as ExtraTrees, XGB, LGBM, RF, Bagging, SVC and KNN, we observed high accuracy greater than 97 % for all the ML models implemented. Upon removing the spatial component (x, y, z) from the OpenFOAM data set used to train / test the ML model to match the accessible data set that can be extracted from FLOW-3D, we high accuracy ranging from 95 % to 97 % were achieved across the investigated model expect KNN which has the lowest value of 90 %. The high accuracy result of ML predictions underscores the efficacy of our ML models in exploring the model for the reduced-order task of spatter prediction, by rigorously testing datasets from the commercial software. The regions associated with spatter in the FLOW-3D simulation output were predicted and visualized on a voxel-by-voxel basis. LPBF melt pool simulations are significantly faster on FLOW-3D than OpenFOAM, LPBF melt pool simulations are significantly faster on FLOW-3D than on OpenFOAM. This makes the screening of a large number of process parameters collected from FLOW-3D feasible. Using an ML model trained on the OpenFOAM dataset, we can develop a process map that displays spatter trends in a large process parameter space. Correlation of the possible amount of spatter with changing power/velocity combination was generated. It was observed that at a high power and low laser speed, more spatters were generated in comparison to low power and high speed.
@@
This research was sponsored by the Army Research Laboratory and conducted under Cooperative Agreement Number W911NF-20-2-0175. The views and conclusions presented in this document are those of the authors and do not necessarily represent the official policies or positions of the Army Research Laboratory or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes, despite any copyright notice herein.
@@
Appendix A Machine Learning Models
A.0.1 Random Forest Classifier (RFClassifier)
Random Forest Classifier is an ensemble learning method that builds numerous decision trees during training and outputs the class that is the most frequent among the individual trees’ predictions [71]. Random forests help mitigate the tendency of the decision trees to overfit their training data [72].
A.0.2 BaggingClassifier
BaggingClassifier is a meta-estimator ensemble learning method that trains base classifiers on random subsets of the original dataset and then combines their individual predictions to produce a final prediction. [73]. This approach helps to minimize the model’s variance as it combines multiple learners’ strengths and smooths out their predictions.
A.0.3 Support Vector Classifier (SVC)
Support Vector Classifier (SVC) is a powerful, flexible, and robust model for classification tasks [74]. It introduced the concept of decision planes to establish decision boundaries. SVC seeks to identify a decision plane that optimizes the margin among different classes in the training set, making it particularly well-suited for binary classification problems [75].
A.1 Ridge Linear Regression (Ridge)
Ridge regression is a regularized version of linear regression, used to reduce overfitting which can occur when the model is over-specified for a small dataset [76]. This method performs $L_{2}$ regularization. Ridge regression shrinks the parameters by imposing a constraint on their values, helping to reduce the model complexity and overfitting.
A.2 The Extra Trees Classifier (Extremely Randomized Trees)
The Extra Trees Classifier (Extremely Randomized Trees) is an ensemble learning method fundamentally similar to Random Forests [67]. However, it differs in the way it constructs the trees. In Extra Trees, randomness is taken a step further in how the splits are determined. In the tree learner, the top-down splitting process is randomized. Instead of calculating the most effective thresholds, thresholds are selected randomly, and the best from the randomly chosen thresholds is used for the splitting rule [77]. This resulted in a reduction in variance, with a trade-off of a slight increase in the model’s bias [78].
A.3 Extreme Gradient Boosting classifier (XGBClassifier)
EXtreme Gradient Boosting classifier (XGBClassifier) implements gradient boosted decision trees designed for speed and performance [79][80]. It is known for its efficiency at handling large datasets and its effectiveness in improving model performance through boosting. XGBoost also includes regularization to reduce overfitting.
A.4 Light Gradient Boosting Machine (LGBMClassifier)
Light Gradient Boosting Machine (LGBMClassifier) is a gradient boosting framework that utilizes tree-based learning algorithms. It is developed for distributed and efficient training, particularly on large datasets [81]. LGBM is known for its high efficiency, handling large data, and lower memory usage. Its efficiency comes from histogram-based optimizations in the tree learning process. LGBMClassifier utilizes tree-based learning algorithms because they offer a balance of accuracy, efficiency, and flexibility, making them highly suitable for structured data and large-scale classification problems. The combination of gradient boosting and optimized tree construction techniques further enhances its performance and scalability.
Figure 16: Trend in power and velocity variation on the volume indicated as spatter.
Figure 17: Relationship between the energy density on the volume indicated as spatter.